# C: A Coverage Path Planning Algorithm for Unknown Environments using Rapidly Covering Graphs

### v1

@@ Abstract
The paper presents a novel sample-based algorithm, called C, for real-time coverage path planning (CPP) of unknown environments. The C algorithm is built upon the concept of a Rapidly Covering Graph (RCG), which tracks the coverage progress and generates the coverage trajectory in uncovered areas. The RCG is constructed incrementally via progressive sampling of the search space as the environment is discovered during robot navigation. The RCG has a sparse-graph structure formed by efficient sampling and pruning techniques; its nodes and edges form the potential waypoints and segments of the robot’s coverage trajectory, respectively. Due to efficient sparse-structure, the RCG produces non-myopic intermediate waypoints of the coverage trajectory. It also guides the robot to escape from the dead-end situations via retreating to the nearest unexplored nodes. A distinct feature of C is that of progressive sampling which eliminates the need for cellular decomposition of the search space. While C produces the desired back-and-forth coverage pattern, it adapts to the TSP-based locally optimal coverage of small uncovered regions, called coverage holes, that are surrounded by obstacles and covered regions. Thus, C proactively detects and covers the coverage holes in situ, which reduces the coverage time by preventing the longer return trajectories from distant regions to cover such holes later.
The algorithmic simplicity and low computational complexity of C makes it easy to implement and suitable for real-time on-board applications. It is analytically proven that C provides complete coverage of unknown environments. The performance of C is validated by 1) extensive high-fidelity simulations and 2) real laboratory experiments using autonomous robots. A comparative evaluation with seven existing CPP methods demonstrate that C yields significant performance improvements in terms of coverage time, number of turns, trajectory length and overlap ratio, while preventing the formation of coverage holes. Finally, C is evaluated on two different applications of CPP using 1) energy-constrained robots and 2) multi-robot teams. A comparative evaluation of C with existing methods on these two applications show similar trends where C trajectories are shorter and with significantly less coverage times.
@@ I Introduction
Recent advancements in sensing, computing and communication technologies have spurred a rapid growth of autonomous service robots that support humans by performing a variety of tasks to improve their quality of life. Specifically, some of the tasks of these robots require covering of all points in the search space. This problem is known as Coverage Path Planning (CPP) [1, 2, 3, 4]. A wide range of application domains exist for CPP including exploration (e.g., mapping of underwater terrains [5, 6]); industry (e.g., structural inspection [7, 8, 9, 10], spray-painting [11, 12] and surface cleaning of vehicles [13]); agriculture (e.g., robotic weeding [14] and arable farming [15]); household (e.g., floor cleaning [16] and lawn mowing [17, 18]); hazard removal (e.g., oil spill cleaning [19, 20]); and defense (e.g., demining [21]).
Figure 1: C-coverage of an island scenario by an AUV.
Often, the environments where coverage operations are performed are unknown. Therefore, it is essential to develop sensor-based CPP methods that can replan and adapt the coverage trajectory in real-time based on dynamic discovery of the environment. It is critical that the CPP method provides complete coverage and does not get stuck in a dead-end. Furthermore, it is desired that the CPP method is efficient, i.e., it minimizes the coverage time by reducing a) the overlaps which in turn reduce the trajectory length, and b) the number of turns. Moreover, it is desired that the coverage method does not create intermediate coverage holes during navigation, which are local uncovered regions surrounded by obstacles and covered regions. This is required to prevent unnecessary inconvenience to cover such holes later via long return trajectories. Finally, it is important that the CPP method is simple to implement on real robot platforms and also computationally efficient to generate adaptive coverage trajectories in real-time.
@@
PPCPP[13]
$\varepsilon^{*}$[3]
BA[22]
NNCPP[23]
BM[24]
BSA[25]
FSSTC[26]
@@
Back and forth
Back and forth
Back and forth
@@
In this regard, the paper presents a novel algorithm, called C, for real-time CPP of unknown environments. Fig. 1 shows an example of C-coverage of an island scenario in an ocean environment by an autonomous underwater vehicle (AUV). While the existing CPP methods rely on different techniques such as cellular decomposition [2], area tiling [3], spanning trees [27], and neural networks [23], C is a sample-based algorithm built upon a novel concept of rapidly covering graphs (RCGs). An RCG is a compact graphical representation of the obstacle-free space, which is incrementally constructed via progressive sampling during robot navigation. The nodes and edges of RCG form the potential waypoints and segments of the robot’s coverage trajectory, respectively. The objectives of RCG are to track the coverage progress, select non-myopic waypoints for the coverage trajectory in uncovered areas and to guide the robot in escaping from the dead-end situations.
C provides complete coverage guarantee of unknown environments using the desired back and forth coverage pattern. During navigation, if the robot detects any coverage holes, which are uncovered regions surrounded by obstacles and covered regions, then C synergistically adapts to the local optimal coverage of such coverage holes by the trajectory generated in situ using the traveling salesman problem (TSP). This eliminates the need to cover such holes later by the longer return trajectories from distant regions. In comparative evaluation with the existing CPP methods, it is shown that C is very efficient and yields significant performance improvements in terms of coverage time, number of turns, path length and overlap ratio. Furthermore, C is simple to implement on real robots and computationally efficient for real-time operations.
@@ I-A Related Work
A review of existing CPP methods is presented in [4]. In general, the CPP methods are classified as random or systematic. While the random methods are simple to implement, they can generate strongly overlapping trajectories and do not guarantee complete coverage in finite time. In contrast, the systematic methods use some basic underlying rules to generate well-defined coverage patterns (e.g., back & forth and spiral) and provide complete coverage in finite time.
@@
The CPP methods are also classified based on the dimension of coverage space. While the 2D methods plan coverage paths to cover the 2D surfaces (e.g., floor cleaning), the 3D methods work in the 3D spaces (e.g., mapping of high-rise buildings by UAVs and underwater terrains by UUVs). Hert et. al. [28] presented a CPP method for underwater terrain mapping using an AUV by sweeping across the terrain while moving at a fixed distance to the uneven surface. Shen and Gupta  [6][29] presented an algorithm, called CT-CPP, that performs layered sensing of the environment using coverage trees (CT) for terrain mapping.
@@
Finally, the CPP methods are classified as offline or online depending on the availability of the environmental information. The offline methods rely on prior information known about the environment to design the coverage paths before the robotic system is deployed for operation. Although these algorithms perform well in known scenarios, their performance can degrade if the prior information is incomplete and/or incorrect. In contrast, the online methods are adaptive [30], i.e., they update the coverage paths in real-time based on the environmental information collected by the onboard sensors.
@@
Several CPP methods have been developed for known environments. Zelinsky et al. [31] proposed a grid-based method, which assigns each grid cell a distance-to-goal value. The coverage path is then generated along the steepest ascent by selecting the unvisited cell in the local neighborhood with the largest distance-to-goal value. Huang [32] developed a CPP method that divided the coverage area into multiple subareas which are each covered by back-and-forth motion with different sweep directions to minimize the total number of turns. Xie et al. [33] proposed a TSP-based method to cover multiple disjoint areas for UAVs. Ramesh et al. [34] computed the optimal partitioning of a non-convex environment for minimizing the number of turns by solving a linear programming problem. Shen et al. [35] presented a neural network based method, called Coverage Path Planning Network (CPPNet), which takes the occupancy grid map of the environment as input and generates a near-optimal coverage path offline.
@@
Several CPP methods have been developed for unknown environments. Lee et al. [36] proposed a grid-based method using a path linking strategy called coarse-to-fine constrained inverse distance transform (CFCIDT) to generate a smooth coverage path, so as to avoid sharp turns. Gabriely and Rimon [27] developed the Spanning Tree Covering (STC) algorithm to generate a systematic spiral path by following a spanning tree. Later in [26], this method was improved to full spiral STC (FS-STC) algorithm. Gonzalez et al. [25] presented the Backtracking Spiral Algorithm (BSA), which generates a spiral path online by following the covered area or obstacles on a specific lateral side until no uncovered free cell can be found in the local neighborhood. Then, a backtracking mechanism is used for escaping from the dead-end situation.
@@
Acar and Choset [2] developed a CPP method which detects the critical points on obstacles to decompose the coverage area into multiple subareas. Then, these subareas and their connections are represented by a graph and coverage is achieved by covering each subarea individually. A limitation of this method is that it can not handle rectilinear environments. This method was further improved by Garcia and Santos  [37]. Ferranti et al. [24] presented the Brick-and-Mortar (BM) algorithm, which gradually enlarges the blocks of inaccessible cells (i.e., visited or obstacle cells) by selecting the target cell as the unexplored cell in the local neighborhood with the most surrounding inaccessible cells. The target cell is marked as visited if it does not block the path between any two accessible cells (i.e., explored or unexplored cells) in the neighborhood; otherwise, it is marked as explored. If there is no unexplored cell in the neighborhood, then an explored cell is selected to traverse through and escape from the dead-end situation.
@@
Luo and Yang [23] developed the Neural Network CPP (NNCPP) algorithm, which places a neuron at the center of every grid cell. The robot path is planned online based on the dynamic activity landscape of the neural network. The activity landscape is updated such that the obstacle cells have negative potential values for collision avoidance and uncovered cells have positive potential values for coverage. The algorithm can escape from the dead-ends by means of neural activity propagation from the unclean areas to attract the robot. Gupta et al. [38] proposed an Ising model approach for dynamic CPP. Viet et al. [22] presented the Boustrophedon motions and A (BA) algorithm, which performs the back-and-forth motion until no uncovered cell is found in the local neighborhood. Then, to escape from a dead-end situation, the algorithm searches for and computes the shortest path to the nearest backtracking cell using the A algorithm [39].
@@
Hassan and Liu [13] proposed the Predator–Prey CPP (PPCPP) algorithm, which is based on a reward function consisting of three factors: 1) predation avoidance reward that maximizes the distance from the predator to prey (predator is a virtual stationary point and prey is the coverage spot of the robot), 2) smoothness reward that minimizes the turn angle for continuing motion in a straight line, and 3) boundary reward for covering boundary targets. Song and Gupta [3] developed the $\varepsilon^{*}$ algorithm, which utilizes an Exploratory Turing Machine (ETM) as a supervisor to guide the robot online with navigation and tasking commands for coverage via back-and-forth motion. To escape from a dead-end, the algorithm uses multi-resolution adaptive potential surfaces (MAPS) to find the unexplored regions.
@@
Recently, some CPP methods have been proposed considering the energy constraints of robots. Shen et al. [40] proposed the $\varepsilon^{*}+$ algorithm which is an extension of the $\varepsilon^{*}$ algorithm [3] for energy-constrained vehicles. It generates the back-and-forth coverage path while continuously monitoring the remaining energy of the robot if it is sufficient to return back to the charging station. Upon battery depletion, the robot retreats back to the charging station using the A algorithm and after recharging it advances to the nearest unexplored cell to restart the coverage of the remaining uncovered area. Shnaps and Rimon [41] developed the battery powered coverage (BPC) algorithm, where the robot follows circular arcs formed by equidistant points to the source. When the battery discharges to a level just sufficient to return to the base, the robot returns to the charging station along the shortest path. Dogru and Marques [42] proposed energy constrained online path planning (ECOCPP) algorithm, which generates coverage and retreat paths by following the contours of covered area.
@@
Some online CPP methods have also been proposed for multi-robot systems [43, 44, 45, 46] and curvature-constrained robots [47, 48, 14, 49].
@@ I-B Summary of the C Algorithm
Fig. 2 illustrates the C operation. While traversing towards a waypoint on the coverage trajectory, the robot senses and discovers the environment. It then incrementally grows the RCG by progressive sampling in the obstacle-free region detected by onboard sensors (details in Section III-B1). The progressive sampling is done in a systematic manner within the sampling front using the frontier samples that are adjacent to the unknown area or obstacles (details in Section III-B2). Then, the nodes of RCG are formed by these samples and serve as the waypoints for the robot. On the other hand, the edges are formed by connecting the nodes and serve as the potential traversal paths for the robot to form the coverage trajectory. The RCG forms a compact graphical representation of the observed obstacle-free space (details in Section III-B3). It has a sparse structure built by efficient sampling and pruning strategies, which provides computational and memory-efficient operation and non-myopic waypoint generation. To track the coverage progress, the nodes of RCG maintain the state as closed (visited) or open (unvisited). The next navigation waypoint of the robot is selected as an unvisited neighboring node on the RCG. The waypoint is selected in a manner to generate the desired back-and-forth coverage pattern (details in Section III-B4). Once the waypoint is selected, the robot moves towards it, maps the environment and covers the region around the trajectory, and the above process is repeated.

### v2

@@ Abstract
The paper presents a novel sample-based algorithm, called C, for real-time coverage path planning (CPP) of unknown environments. C is built upon the concept of a Rapidly Covering Graph (RCG), which is incrementally constructed during robot navigation via progressive sampling of the search space. By using efficient sampling and pruning techniques, the RCG is constructed to be a minimum-sufficient graph, where its nodes and edges form the potential waypoints and segments of the coverage trajectory, respectively. The RCG tracks the coverage progress, generates the coverage trajectory and helps the robot to escape from the dead-end situations. To minimize coverage time, C produces the desired back-and-forth coverage pattern, while adapting to the TSP-based optimal coverage of local isolated regions, called coverage holes, which are surrounded by obstacles and covered regions. It is analytically proven that C provides complete coverage of unknown environments. The algorithmic simplicity and low computational complexity of C make it easy to implement and suitable for real-time on-board applications. The performance of C is validated by 1) extensive high-fidelity simulations and 2) laboratory experiments using an autonomous robot. C yields near optimal trajectories, and a comparative evaluation with seven existing CPP methods demonstrates significant improvements in performance in terms of coverage time, number of turns, trajectory length, and overlap ratio, while preventing the formation of coverage holes. Finally, C is comparatively evaluated on two different CPP applications using 1) energy-constrained robots and 2) multi-robot teams.
@@ I Introduction
Recent advancements in sensing, computing and communication technologies have spurred a rapid growth of autonomous service robots that support humans by performing a variety of tasks to improve their quality of life. Specifically, some of these tasks require planning a path to cover all points in the search space, this problem is called CPP [acar2006sensor, acar2002sensor, song2018, galceran2013survey]. A wide range of application domains exist for CPP including exploration (e.g., mapping of underwater terrains [palomeras2018, shen2022ct]); industry (e.g., structural inspection [englot2013three, vidal2017, bircher2018receding, song2020online], spray-painting [vempati2018, atkar2005] and surface cleaning of vehicles [hassan2019ppcpp]); agriculture (e.g., robotic weeding [maini2022online] and arable farming [jin2011]); household (e.g., floor cleaning [palacin2004building] and lawn mowing [weiss2008lawn, Song_sose2015]); hazard removal (e.g., oil spill cleaning [SGH13]); and defense (e.g., demining [mukherjee2011symbolic]).
Often, the environments where coverage operations are performed are unknown. Therefore, it is essential to develop sensor-based CPP methods that can replan and adapt the coverage trajectory in real-time based on dynamic discovery of the environment. It is critical that the CPP method provides complete coverage and does not get stuck in a dead-end. Furthermore, it is desired that the CPP method is efficient, i.e., it minimizes the coverage time by reducing a) the overlaps which in turn reduces the trajectory length, and b) the number of turns. Moreover, it is desired that the coverage method does not create intermediate coverage holes during navigation, which are local uncovered regions surrounded by obstacles and covered regions. This is required to prevent unnecessary inconvenience to cover such holes later via long return trajectories. Finally, it is important that the CPP method is simple to implement on real robot platforms and also computationally efficient to generate adaptive coverage trajectories in real-time.
Figure 1: C coverage of an island scenario by an AUV.
@@
PPCPP[hassan2019ppcpp]
$\varepsilon^{*}$[song2018]
BA[viet2013ba]
NNCPP[luo2008bioinspired]
BM[ferranti2007brick]
BSA[gonzalez2005bsa]
FSSTC[gabriely2003competitive]
@@
Back-and-forth
Back-and-forth
Back-and-forth
@@
In this regard, the paper presents a novel algorithm, called C, for real-time CPP of unknown environments. Fig. 1 shows an example of C-coverage of an island scenario in an ocean environment by an autonomous underwater vehicle (AUV). While the existing CPP methods rely on different techniques such as cellular decomposition [acar2002sensor], area tiling [song2018], spanning trees [gabriely2001spanning], and neural networks [luo2008bioinspired], C is a sample-based algorithm built upon a novel concept of RCGs. An RCG is a minimum sufficient graphical representation of the obstacle-free space, which is incrementally constructed via progressive sampling during robot navigation. The nodes and edges of an RCG form the potential waypoints and segments of the robot’s coverage trajectory, respectively. The objectives of RCG are to track the coverage progress, select non-myopic waypoints for the coverage trajectory in uncovered areas, and guide the robot to escape from dead-end situations.
C provides a complete coverage guarantee of unknown environments using the desired back-and-forth coverage pattern. During navigation, if the robot detects the formation of any potential coverage holes, then C synergistically adapts to the locally optimal trajectories generated in situ using the traveling salesman problem (TSP) to cover such regions. This prevents the formation of undesired coverage holes in the map during the coverage process, thus eliminating the need to cover them later by return trajectories from distant regions. C is validated by high-fidelity simulations and real experiments, which demonstrate that C is highly efficient and produces near optimal trajectories. A comparative evaluation with the existing CPP methods shows that C yields significant performance improvements in terms of coverage time, number of turns, trajectory length and overlap ratio. Furthermore, C is simple to implement and computationally efficient to run on real robotic platforms for real-time operations.
@@ I-A Related Work
A review of existing CPP methods is presented in [galceran2013survey]. CPP methods can be classified into various categories.
I-A1 Random vs Systematic Methods
@@
In general, the CPP methods are classified as random or systematic. While the random methods are simple to implement, they can generate strongly overlapping trajectories and do not guarantee complete coverage in finite time. In contrast, the systematic methods use some basic underlying rules to generate well-defined coverage patterns (e.g., back & forth and spiral) and provide complete coverage in finite time.
I-A2 2D vs 3D Methods
@@
The CPP methods are also classified based on the dimension of coverage space. While the 2D methods plan coverage paths to cover the 2D surfaces (e.g., floor cleaning), the 3D methods work in the 3D spaces (e.g., mapping of high-rise buildings by UAVs and underwater terrains by UUVs). Hert et. al. [hert1996] presented a CPP method for underwater terrain mapping using an AUV by sweeping across the terrain while moving at a fixed distance to the uneven surface. Shen and Gupta [shen2022ct] presented an algorithm, called CT-CPP, that performs layered sensing of the environment using coverage trees (CT) for terrain mapping.
I-A3 Offline vs Online Methods
@@
The CPP methods are also classified as offline or online depending on the availability of the environmental information. The offline methods rely on prior information known about the environment to design the coverage paths before the robotic system is deployed for operation. Although these algorithms perform well in known scenarios, their performance can degrade if the prior information is incomplete and/or incorrect. In contrast, the online methods are adaptive [Shen_SMART2023], i.e., they update the coverage paths in real-time based on the environmental information collected by the onboard sensors.
@@
$\bullet$ Known environments: Several CPP methods have been developed for known environments. Zelinsky et al. [zelinsky1993planning] proposed a grid-based method, which assigns each grid cell a distance-to-goal value. The coverage path is then generated along the steepest ascent by selecting the unvisited cell in the local neighborhood with the largest distance-to-goal value. Huang [huang2001optimal] developed a CPP method that divided the coverage area into multiple subareas which are each covered by back-and-forth motion with different sweep directions to minimize the total number of turns. Xie et al. [xie2020path] proposed a TSP-based method to cover multiple disjoint areas for UAVs. Ramesh et al. [ramesh2022optimal] computed the optimal partitioning of a non-convex environment for minimizing the number of turns by solving a linear programming problem. Shen et al. [shen2021cppnet] presented a neural network based method, called CPP Network (CPPNet), which takes the occupancy grid map of the environment as input and generates a near-optimal coverage path offline.
@@
$\bullet$ Unknown environments: Several CPP methods have been developed for unknown environments. Lee et al. [LBCO11] proposed a grid-based method using a path linking strategy called coarse-to-fine constrained inverse distance transform (CFCIDT) to generate a smooth coverage path, so as to avoid sharp turns. Gabriely and Rimon [gabriely2001spanning] developed the Spanning Tree Covering (STC) algorithm, which discretizes the workspace into coarse cells, achieving complete coverage with zero overlap but at a coarser resolution than the robot’s footprint, which may leave gaps. Full spiral STC (FS-STC) [gabriely2003competitive] extends STC by incorporating partially occupied cells into an augmented graph, enabling fine-cell level coverage at the resolution of the robot’s footprint. Gonzalez et al. [gonzalez2005bsa] presented the Backtracking Spiral Algorithm (BSA), which generates a spiral path online by following the covered area or obstacles on a specific lateral side until no uncovered free cell can be found in the local neighborhood. Then, a backtracking mechanism is used to escape from a dead-end.
@@
Acar and Choset [acar2002sensor] developed a CPP method which detects the critical points on obstacles to decompose the coverage area into multiple subareas. Then, these subareas and their connections are represented by a graph and coverage is achieved by covering each subarea individually. A limitation of this method is that it can not handle rectilinear environments. This method was further improved by Garcia and Santos [garcia2004mobile]. Ferranti et al. [ferranti2007brick] presented the Brick-and-Mortar (BM) algorithm, which gradually enlarges the blocks of inaccessible cells (i.e., visited or obstacle cells) by selecting the target cell as the unexplored cell in the local neighborhood with the most surrounding inaccessible cells. The target cell is marked as visited if it does not block the path between any two accessible cells (i.e., explored or unexplored cells) in the neighborhood; otherwise, it is marked as explored. If there is no unexplored cell in the neighborhood, then an explored cell is selected to traverse through and escape from the dead-end situation.
@@
Luo and Yang [luo2008bioinspired] developed the Neural Network CPP (NNCPP) algorithm, which places a neuron at the center of every grid cell. The robot path is planned online based on the dynamic activity landscape of the neural network. The activity landscape is updated such that the obstacle cells have negative potential values for collision avoidance and uncovered cells have positive potential values for coverage. The algorithm can escape from the dead-ends by means of neural activity propagation from the unclean areas to attract the robot. Gupta et al. [GRP09] proposed an Ising model approach for dynamic CPP. Viet et al. [viet2013ba] presented the Boustrophedon motions and A (BA) algorithm, which performs the back-and-forth motion until no uncovered cell is found in the local neighborhood. Then, to escape from a dead-end situation, the algorithm searches for and computes the shortest path to the nearest backtracking cell using the A algorithm [hart1968formal].
@@
Hassan and Liu [hassan2019ppcpp] proposed the Predator–Prey CPP (PPCPP) algorithm, which is based on a reward function consisting of three factors: 1) predation avoidance reward that maximizes the distance from the predator to prey (predator is a virtual stationary point and prey is the coverage spot of the robot), 2) smoothness reward that minimizes the turn angle for continuing motion in a straight line, and 3) boundary reward for covering boundary targets. Song and Gupta [song2018] developed the $\varepsilon^{*}$ algorithm, which utilizes an Exploratory Turing Machine (ETM) as a supervisor to guide the robot online with navigation and tasking commands for coverage via back-and-forth motion. To escape from a dead-end, the algorithm uses multi-resolution adaptive potential surfaces (MAPS) to find the unexplored regions.
I-A4 Energy Constrained Methods
@@
Recently, some CPP methods have been proposed considering the energy constraints of robots. Shen et al. [shen2020] proposed the $\varepsilon^{*}+$ algorithm which is an extension of the $\varepsilon^{*}$ algorithm [song2018] for energy-constrained vehicles. It generates the back-and-forth coverage path while continuously monitoring the remaining energy of the robot if it is sufficient to return back to the charging station. Upon battery depletion, the robot retreats back to the charging station using the A algorithm and after recharging it advances to the nearest unexplored cell to restart the coverage of the remaining uncovered area. Shnaps and Rimon [shnaps2016] developed the battery powered coverage (BPC) algorithm, where the robot follows circular arcs formed by equidistant points to the source. When the battery discharges to a level just sufficient to return to the base, the robot returns to the charging station along the shortest path. Dogru and Marques [dogru2022eco] proposed energy constrained online path planning (ECOCPP) algorithm, which generates coverage and retreat paths by following the contours of covered area.
I-A5 Multi-robot Methods
@@
Some online CPP methods have also been proposed for multi-robot systems [palacios2019equitable, song2020care, hassan2020dec, wang2024apf] and curvature-constrained robots [shen2019online, kan2020online, maini2022online, GMDM_TRO].
@@ I-B Summary of the C Algorithm
Fig. 2 illustrates the C operation. While traversing towards a waypoint on the coverage trajectory, the robot senses and discovers the environment. It then incrementally grows the RCG by progressive sampling in the obstacle-free region detected by onboard sensors (details in Section III-B1). The progressive sampling is done in a systematic manner within the sampling front using the frontier samples that are adjacent to the unknown area or obstacles (details in Section III-B2). Then, the nodes of RCG are formed by these samples and serve as the waypoints for the robot. On the other hand, the edges are formed by connecting the nodes and serve as the potential traversal paths for the robot to form the coverage trajectory. The RCG forms a minimum sufficient graphical representation of the observed obstacle-free space (details in Section III-B3). It has a sparse structure built by efficient sampling and pruning strategies, which provides computational and memory-efficient operation and non-myopic waypoint generation.
To track the coverage progress, the nodes of RCG maintain the state as closed (visited) or open (unvisited). The next waypoint of the robot is selected as an unvisited neighboring node on the RCG. The waypoint is selected in a manner to generate the desired back-and-forth coverage pattern (details in Section III-B4). Once the waypoint is selected, the robot moves towards it, maps the environment, and covers the region around the trajectory. The above process is then repeated.
C has two important features, as described below.